# Beyond Linear Approximations: A Novel Pruning Approach for Attention Matrix

### v1

@@ 5.1 Previous Tools on Convergence of GD
We say that a function $f:\mathbb{R}^{p}\to\mathbb{R}$ satisfies a $g$-proxy, $\xi$-optimal Polyak–Łojasiewicz inequality with parameters $\alpha>0$ and $\mu>0$ (in short, $f$satisfies the $(g,\xi,\alpha,\mu)$-PL inequality) if there exists a function $g:\mathbb{R}^{p}\to\mathbb{R}$ and scalars $\xi\in\mathbb{R}$, $\mu>0$ such that for all $w\in\mathbb{R}^{p}$,
$\|\nabla f(w)\|^{\alpha}\geq\frac{1}{2}\mu(g(w)-\xi).$
@@
Suppose $f(w)$ satisfies the $(g(\cdot),\xi,\alpha,\mu)$-proxy PL inequality for some function $g(\cdot):\mathbb{R}^{p}\to\mathbb{R}$. Assume that $f$ is non-negative and has $L$-Lipschitz gradients. Then for any $\epsilon>0$, provided $\eta<1/L$, GD with fixed step size $\eta$ and run for $T=2\eta^{-1}(\mu\epsilon/2)^{-2/\alpha}f(w^{(0)})$ iterations results in the following guarantee,
$\min_{t<T}g(w^{(t)})\leq\xi+\epsilon.$
@@ 5.2 Closed Form of Gradient
As a first step, we compute the close form of the gradient $\nabla_{M}{\cal L}(M)$. The pruning mask $M$ is inside a non-linear function Softmax, which complicates our calculation. We defer the proof to Section B.
Theorem 5.3 (Closed form of gradient, informal version of Theorem C.5).
Let ${\cal L}(M)$ be defined in Definition 3.2. Let $p$ be defined in Definition C.1. Let $X\in\mathbb{R}^{n\times d}$, $M\in[0,1]^{d\times d}$, $W\in\mathbb{R}^{d\times d}$. Then, we have
@@ 5.3 Lipschitz of Gradient
Theorem 5.4 (Lipschitz of the gradient, informal version of Theorem E.8).
Let $R$ be some fixed constant satisfies $R>1$. Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$. We have $\|X\|_{F}\leq R$ and $\|W\|_{F}\leq R$. Let ${\cal L}(M)$ be defined in Definition 3.2. For $M,\widetilde{M}\in\mathbb{R}^{d\times d}$, we have
@@
We defer the proof to Section E. Establishing the Lipschitz continuity of the gradient satisfies one of the necessary conditions for applying Theorem 5.2. The above theorem implicates that the gradient for $M$ is upper bounded, providing a way to choose step size.
@@ 5.4 PL Inequality of Gradient
Lemma 5.5 (PL inequality, informal version of Lemma F.10).
@@
We have
$\|\nabla_{M}\mathcal{L}(M)\|_{F}^{2}\geq\frac{1}{2}\mu(2\mathcal{L}_{\mathrm{attn}}(M)+\frac{2\lambda^{2}}{\mu}\|M\|_{F}^{2}-\xi).$
We defer the proof to Section F. By confirming that ${\cal L}(M)$ satisfies the PL inequality and that its gradient is Lipschitz continuous, we then apply Theorem 5.2 to conclude that GD will converge to a solution within our desired error tolerance, and further prove Theorem 4.1.
@@
Lemma 5.6 (Informal version of Lemma F.4).
Let $B\in\mathbb{R}^{n\times n}$ and $X\in\mathbb{R}^{n\times d}$. Assume that $XX^{\top}\succeq\beta I$. Then, we have
$\|X^{\top}BX\|_{F}\geq\beta\|B\|_{F}.$
Lemma 5.7 (Informal version of Lemma F.7).
Let $B\in\mathbb{R}^{n\times n}$ and each row summation is zero, i.e., $B\cdot{\bf 1}_{n}={\bf 0}_{n}$. Let $\widetilde{B}\in[0,1]^{n\times n}$ and each row summation is 1, i.e., $\widetilde{B}\cdot{\bf 1}_{n}={\bf 1}_{n}$. Assume that $\min_{i,j\in[n]}\widetilde{B}_{i,j}\geq\delta>0$. Then, we can show
$\|B\circ\widetilde{B}-\operatorname{diag}((B\circ\widetilde{B})\cdot{\bf 1}_{n})\widetilde{B}\|_{F}\geq\delta\cdot\|B\|_{F}.$
@@ 6 Experiment
In this section, we discuss the experiments conducted to illustrate the effectiveness of our Algorithm 1. We first introduce our settings in Section 6.1. Then, we present our results in Section 6.2.
6.1 Settings
Method and evaluation.
@@ 6.1 Settings
We implement our method following the pseudocode in Algorithm 1, using NumPy and JAX for acceleration. We evaluate our method on unstructured sparsity, meaning that zeros can occur anywhere within the attention weight matrix $W$. Specifically, we use Definition 3.2 as our loss function, optimizing over the pruning mask $M$ using gradient descent based on the closed-form expression derived in Theorem 5.3. To accelerate convergence, we leverage momentum into the optimization process and fix the momentum parameter at $0.9$. After obtaining the optimal pruning mask, we convert $M$ to a binary pruning mask to prune $W$, maintaining sparsity at the desired pruning ratio $\rho$. We use the relative error as our evaluation metric, which is defined as
$\frac{\|\widetilde{D}^{-1}\widetilde{A}-D^{-1}A\|_{F}^{2}}{\|D^{-1}A\|_{F}^{2}},$
where $\widetilde{D}$, $\widetilde{A}$, $D$, $A$ are defined in Definition 3.2.
Baselines.
@@
We compare our method with two linear pruning approaches, namely Wanda [75] and SparseGPT [26]. Wanda is a pruning method that removes weights with the smallest magnitudes multiplied by the corresponding input activations, achieving sparsity without requiring retraining or weight updates. SparseGPT is a second-order pruning method that utilizes the Hessian matrix to prune a portion of the weight matrix while simultaneously updating the remaining parameters. We implement Wanda and SparseGPT as described in their respective papers. Notably, since the settings of SparseGPT and Wanda are linear, we do not prune the fused weight matrix $W$ directly; instead, we prune $W_{Q}$ and $W_{K}$ separately (see Figure 1).
Data.
@@
In order to assess the efficacy of different methods in approximating the attention matrix, we construct the data via a carefully defined generating process. Specifically, we create multiple independent random Gaussian matrices $G\in\mathbb{R}^{d\times d}$, where each entry of $G$ drawn from a normal distribution i.e., $G_{i,j}\sim\mathcal{N}(0,1)$ for $i,j\in[d]$. Then, we perform singular value decomposition (SVD) on matrix $G$, i.e., $U,S,V^{\top}=\mathrm{SVD}(G)$. We retain the first four singular values in $S$ and set others to zero, constraining the rank to four. Our $W_{Q}$ and $W_{K}$ are then constructed as $U\operatorname{diag}(S)V^{\top}$. The weight matrix $W$ used in our setting is formed by taking the product $W=W_{Q}W_{K}^{\top}$. For $X\in\mathbb{R}^{n\times d}$, we generate it as a full-rank Gaussian random matrix.
Setup.
@@
In our experiments, the weight matrix dimension $d=64$ is kept as constant across all figures, and we simulate two datasets of size $k=16$ and $k=64$. We set the input sequence length $n=128$ for experiments (a) and (c) in Figure 2. The pruning ratio $\rho=0.5$ is set for experiments (a) and (b) in Figure 2. For our method, the regularization coefficient $\lambda:=\widetilde{\lambda}/n$ where we abuse the notation to denote $\widetilde{\lambda}$ as the same used in Definition 3.2 and $\lambda$ here is the parameter we really control in experiments. $\lambda$ is set as $0.04$ for experiments (b) and (c) in Figure 2 (intuition drawn from experiment (a)). The total number of epochs is set as $T=100$. The step size is set as $\eta=0.1/\lambda$ because Theorem 4.1 indicates that $\eta$ is inversely proportional to $\lambda$ with some constant, i.e., $\eta\propto 1/L\propto 1/(\lambda+\mathrm{other~{}terms})$.
6.2 Results
@@
Overall, the results in Figure 2 show that our Algorithm 1 outperforms Wanda and SparseGPT with a large margin, which supports our theoretical analysis in Theorem 4.1. In the following, we will discuss each setting in detail.
Relation with regularization coefficient $\lambda$.
@@
The leftmost column of Figure 2 investigates the impact of the regularization coefficient $\lambda$ on relative error. As $\lambda$ increases from very small values, the relative error initially decreases sharply for our algorithm, reaching a minimum before gradually rising again, which forms a $U$ shape curve. This behavior indicates that there is an optimal $\lambda$ where our algorithm achieves its best performance around $2^{-4}$. The U-shape curve phenomena are well-known in most hyper-parameter choosing, e.g., regularization coefficient.
Relation with input sequence length $n$.
@@
The center column of Figure 2 explores how the relative error changes with respect to the input sequence length $n$. As $n$ increases, the relative error for all three methods grows, though at different rates. Our method demonstrates a slower increase, maintaining a significant margin over both Wanda and SparseGPT, particularly for larger values of $n$. Wanda, while showing better performance than SparseGPT for larger sequence lengths, becomes comparable to SparseGPT as $n$ is relatively small.
Relation with pruning ratio $\rho$.
@@
The rightmost column of Figure 2 illustrates the relationship between the relative error and the pruning ratio $\rho$ for the three methods under comparison: our algorithm, Wanda, and SparseGPT. As the pruning ratio $\rho$ increases, all methods exhibit a rise in relative error, indicating a degradation in approximation accuracy. However, our algorithm consistently outperforms both Wanda and SparseGPT across the range of $\rho$, with a lower relative error. SparseGPT and Wanda follow a similar trend, closely tracking each other.

### v2

@@ 5.1 Previous Tools on Convergence of GD
We say that a function $f:\mathbb{R}^{p}\to\mathbb{R}$ satisfies a $g$-proxy, $\xi$-optimal Polyak–Łojasiewicz inequality with parameters $\alpha>0$ and $\mu>0$ (in short, $f$satisfies the $(g,\xi,\alpha,\mu)$-PL inequality) if there exists a function $g:\mathbb{R}^{p}\to\mathbb{R}$ and scalars $\xi\in\mathbb{R}$, $\mu>0$ such that for all $w\in\mathbb{R}^{p}$, $\|\nabla f(w)\|^{\alpha}\geq\frac{1}{2}\mu(g(w)-\xi).$
@@
Suppose $f(w)$ satisfies the $(g(\cdot),\xi,\alpha,\mu)$-proxy PL inequality for some function $g(\cdot):\mathbb{R}^{p}\to\mathbb{R}$. Assume that $f$ is non-negative and has $L$-Lipschitz gradients. Then for any $\epsilon>0$, provided $\eta<1/L$, GD with fixed step size $\eta$ and run for $T=2\eta^{-1}(\mu\epsilon/2)^{-2/\alpha}f(w^{(0)})$ iterations results in the following guarantee, $\min_{t<T}g(w^{(t)})\leq\xi+\epsilon.$
@@ 5.2 Closed Form of Gradient
As a first step, we compute the close form of the gradient $\nabla_{M}{\cal L}(M)$. The pruning mask $M$ is inside a non-linear function Softmax, which complicates our calculation. We defer the proof to Section C.
Theorem 5.3 (Closed form of gradient, informal version of Theorem D.5).
Let ${\cal L}(M)$ be defined in Definition 3.2. Let $p$ be defined in Definition D.1. Let $X\in\mathbb{R}^{n\times d}$, $M\in[0,1]^{d\times d}$, $W\in\mathbb{R}^{d\times d}$. Then, we have
@@ 5.3 Lipschitz of Gradient
Theorem 5.4 (Lipschitz of the gradient, informal version of Theorem F.8).
Let $R$ be some fixed constant that satisfies $R>1$. Let $X\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{d\times d}$. We have $\|X\|_{F}\leq R$ and $\|W\|_{F}\leq R$. Let ${\cal L}(M)$ be defined in Definition 3.2. For $M,\widetilde{M}\in\mathbb{R}^{d\times d}$, we have
@@
We defer the proof to Section F. Establishing the Lipschitz continuity of the gradient satisfies one of the necessary conditions for applying Theorem 5.2. The above theorem implicates that the gradient for $M$ is upper bounded, providing a way to choose step size.
@@ 5.4 PL Inequality of Gradient
Lemma 5.5 (PL inequality, informal version of Lemma G.10).
@@
We have $\|\nabla_{M}\mathcal{L}(M)\|_{F}^{2}\geq\frac{1}{2}\mu(2\mathcal{L}_{\mathrm{attn}}(M)+\frac{2\lambda^{2}}{\mu}\|M\|_{F}^{2}-\xi).$
We defer the proof to Section G. By confirming that ${\cal L}(M)$ satisfies the PL inequality and that its gradient is Lipschitz continuous, we then apply Theorem 5.2 to conclude that GD will converge to a solution within our desired error tolerance, and further prove Theorem 4.1.
@@
Lemma 5.6 (Informal version of Lemma G.4).
Let $B\in\mathbb{R}^{n\times n}$ and $X\in\mathbb{R}^{n\times d}$. Assume that $XX^{\top}\succeq\beta I$. Then, we have $\|X^{\top}BX\|_{F}\geq\beta\|B\|_{F}.$
Lemma 5.7 (Informal version of Lemma G.7).
Let $B\in\mathbb{R}^{n\times n}$ and each row summation is zero, i.e., $B\cdot{\bf 1}_{n}={\bf 0}_{n}$. Let $\widetilde{B}\in[0,1]^{n\times n}$ and each row summation is 1, i.e., $\widetilde{B}\cdot{\bf 1}_{n}={\bf 1}_{n}$. Assume that $\min_{i,j\in[n]}\widetilde{B}_{i,j}\geq\delta>0$. Then, we can show $\|B\circ\widetilde{B}-\operatorname{diag}((B\circ\widetilde{B})\cdot{\bf 1}_{n})\widetilde{B}\|_{F}\geq\delta\cdot\|B\|_{F}.$
@@ 6 Experiment
6.1 Evaluation on Synthetic Data
We discuss the synthetic experiments conducted to illustrate the effectiveness of our Algorithm 1.
@@ 6.1 Evaluation on Synthetic Data
Method and evaluation. We implement our method following the pseudocode in Algorithm 1, using NumPy and JAX for acceleration. We evaluate our method on unstructured sparsity, meaning that zeros can occur anywhere within the attention weight matrix $W$. Specifically, we use Definition 3.2 as our loss function, optimizing over the pruning mask $M$ using gradient descent based on the closed-form expression derived in Theorem 5.3. To accelerate convergence, we leverage momentum into the optimization process and fix the momentum parameter at $0.9$. After obtaining the optimal pruning mask, we convert $M$ to a binary pruning mask to prune $W$, maintaining sparsity at the desired pruning ratio $\rho$. We use the relative error as our evaluation metric, which is defined as ${\|\widetilde{D}^{-1}\widetilde{A}-D^{-1}A\|_{F}^{2}}/{\|D^{-1}A\|_{F}^{2}},$ where $\widetilde{D}$, $\widetilde{A}$, $D$, $A$ are defined in Definition 3.2.
@@
Baselines. We compare our method with two linear pruning approaches, namely Wanda [120] and SparseGPT [40]. Wanda is a pruning method that removes weights with the smallest magnitudes multiplied by the corresponding input activations, achieving sparsity without requiring retraining or weight updates. SparseGPT is a second-order pruning method that utilizes the Hessian matrix to prune a portion of the weight matrix while simultaneously updating the remaining parameters. We implement Wanda and SparseGPT as described in their respective papers. Notably, since the settings of SparseGPT and Wanda are linear, we do not prune the fused weight matrix $W$ directly; instead, we prune $W_{Q}$ and $W_{K}$ separately (see Figure 1).
@@
Data. In order to assess the efficacy of different methods in approximating the attention matrix, we construct the data via a carefully defined generating process. Specifically, we create multiple independent random Gaussian matrices $G\in\mathbb{R}^{d\times d}$, where each entry of $G$ drawn from a normal distribution, i.e., $G_{i,j}\sim\mathcal{N}(0,1)$ for $i,j\in[d]$. Then, we perform singular value decomposition (SVD) on matrix $G$, i.e., $U,S,V^{\top}=\mathrm{SVD}(G)$. We retain the first four singular values in $S$ and set others to zero, constraining the rank to four. Our $W_{Q}$ and $W_{K}$ are then constructed as $U\operatorname{diag}(S)V^{\top}$. The weight matrix $W$ used in our setting is formed by taking the product $W=W_{Q}W_{K}^{\top}$. For $X\in\mathbb{R}^{n\times d}$, we generate it as a full-rank Gaussian random matrix.
@@
Setup. In our experiments, the weight matrix dimension $d=64$ is kept constant across all figures, and we simulate two datasets of size $k=16$ and $k=64$. We set the input sequence length $n=128$ for experiments (a) and (c) in Figure 2. The pruning ratio $\rho=0.5$ is set for experiments (a) and (b) in Figure 2. For our method, the regularization coefficient $\lambda:=\widetilde{\lambda}/n$ where we abuse the notation to denote $\widetilde{\lambda}$ as the same used in Definition 3.2 and $\lambda$ here is the parameter we really control in experiments. $\lambda$ is set as $0.04$ for experiments (b) and (c) in Figure 2 (intuition drawn from experiment (a)). The total number of epochs is set as $T=100$. The step size is set as $\eta=0.1/\lambda$ because Theorem 4.1 indicates that $\eta$ is inversely proportional to $\lambda$ with some constant, i.e., $\eta\propto 1/L\propto 1/(\lambda+\mathrm{other~{}terms})$.
@@
Results. Overall, the results in Figure 2 show that our Algorithm 1 outperforms Wanda and SparseGPT with a large margin, which supports our theoretical analysis in Theorem 4.1. In the following, we will discuss each setting in detail.
@@
Relation with regularization coefficient $\lambda$. The leftmost column of Figure 2 investigates the impact of the regularization coefficient $\lambda$ on relative error. As $\lambda$ increases from very small values, the relative error initially decreases sharply for our algorithm, reaching a minimum before gradually rising again, which forms a $U$ shape curve. This behavior indicates that there is an optimal $\lambda$ where our algorithm achieves its best performance around $2^{-4}$. The U-shape curve phenomena are well-known in most hyper-parameter choosing, e.g., regularization coefficient.
@@
Relation with input sequence length $n$. The center column of Figure 2 explores how the relative error changes with respect to the input sequence length $n$. As $n$ increases, the relative error for all three methods grows, though at different rates. Our method demonstrates a slower increase, maintaining a significant margin over both Wanda and SparseGPT, particularly for larger values of $n$. Wanda, while showing better performance than SparseGPT for larger sequence lengths, becomes comparable to SparseGPT as $n$ is relatively small.
@@
Relation with pruning ratio $\rho$. The rightmost column of Figure 2 illustrates the relationship between the relative error and the pruning ratio $\rho$ for the three methods under comparison: our algorithm, Wanda, and SparseGPT. As the pruning ratio $\rho$ increases, all methods exhibit a rise in relative error, indicating a degradation in approximation accuracy. However, our algorithm consistently outperforms both Wanda and SparseGPT across the range of $\rho$, with a lower relative error. SparseGPT and Wanda follow a similar trend, closely tracking each other.
Figure 3: The comparison among our algorithm, Wanda, and SparseGPT on Llama 3.2-1B.
6.2 Experiment on Real Dataset and LLMs
In this subsection, we discuss the experiments conducted on the real dataset and LLMs to illustrate the effectiveness of our method.
Method and evaluation. We evaluate our method on unstructured sparsity, meaning that zeros can occur anywhere within the attention weight matrices $W_{Q}$ and $W_{K}$. Specifically, we use the loss function defined below:
${\cal L}(M_{Q},M_{K}):=\frac{1}{2}\|D^{-1}A-\widetilde{D}^{-1}\widetilde{A}\|_{F}^{2}+\frac{1}{2}(\|M_{Q}\|_{F}^{2}+\|M_{K}\|_{F}^{2})$ (1)
where $D$ and $A$ are the original attention matrix, $M_{Q},M_{K}$ are pruning masks, and
$\widetilde{A}:=\exp(X(M_{Q}\circ W_{Q})(M_{K}\circ W_{K})^{\top}X^{\top}),~{}~{}~{}~{}\widetilde{D}:=\operatorname{diag}(\widetilde{A}\cdot\textbf{1}_{n}).$
After obtaining the optimal pruning mask, we convert the pruning mask to a binary pruning mask to prune $W_{Q}$ and $W_{K}$, maintaining sparsity at the desired pruning ratio $\rho$. We use the relative error as our evaluation metric.
Baselines. The pruning is performed on the last attention layer of the pretrained model Llama 3.2-1B [102]. The baselines are Wanda and SparseGPT, the same as Section 6.2.
Data. To simulate real-world large language models, we utilize the Colossal Clean Crawled Corpus (C4 Dataset) [115], which is also used as the calibration dataset in our baselines Wanda and SparseGPT. Additionally, with a primary focus on pruning the attention matrix, we extract the input hidden states corresponding to the target attention matrix from the pretrained Llama 3.2 model using a customized hook function and use these as our input $X$.
Setup. The weight matrix dimension is $d=2048$ in Llama 3.2-1B. For all the experiments, we set $\lambda$ as $0.05$ and $\eta$ as $0.005$. For the varying $n$ experiment, we set the pruning ratio $\rho$ as $0.7$. For the varying $\rho$ experiment, we set the input sequence length $n$ as $512$.
Results. Overall, the results in Figure 3 show that our algorithm outperforms Wanda and SparseGPT in real-world LLMs, which supports our theoretical analysis in Theorem 4.1 and enhances our preliminary experiment in Section 6. In the following, we will discuss each setting in detail.
Relation with input sequence length $n$. The left column of Figure 3 shows that our algorithm outperforms the baselines in different sequence lengths continuously.
Relation with pruning ratio $\rho$. The right column of Figure 3 illustrates that as the pruning ratio $\rho$ increases, all methods exhibit a rise in relative error. But our algorithm consistently outperforms both Wanda and SparseGPT across the range of $\rho$ with a much lower increasing rate.
Assumptions verification. Notice that Theorem 4.1 relies on two assumptions: $XX^{\top}\succeq\beta I$ and $\min_{i,j\in[n]}(\widetilde{D}^{-1}\widetilde{A})_{i,j}\geq\delta>0$. We verify these assumptions using the C4 dataset, obtaining $\beta\approx 0.034$ and $\delta\approx 0.0025$, thereby demonstrating the practicality of our theoretical framework.
Table 1: Comparison of different methods based on Perplexity (PPL).
Method PPL
Dense MLP + Dense Attn 12.487
Dense MLP + SparseGPT Attn 14.269
Dense MLP + Wanda Attn 14.912
Dense MLP + Our Attn 13.885
Wanda MLP + Wanda Attn 30.426
Wanda MLP + SparseGPT Attn 26.074
Wanda MLP + Our Attn 24.427
SparseGPT MLP + Wanda Attn 45.435
SparseGPT MLP + SparseGPT Attn 36.641
SparseGPT MLP + Our Attn 34.946
6.3 Experiment on End-to-end Perplexity
We present the experiment on end-to-end perplexity in this section.
Baselines. As our method focuses on pruning the attention matrix, it can be seamlessly combined with approaches that perform linear pruning on the MLP, such as Wanda and SparseGPT. Therefore, we conduct three groups of experiments: (1) using a dense MLP without pruning, (2) applying SparseGPT to prune the MLP layer, and (3) using Wanda to prune the MLP layer. Subsequently, we apply Wanda, SparseGPT, and our method to prune the attention weights.
Method and data. We use the same method and data as Section 6.2.
Setup. We use Llama 3.2 1B as the target model to prune, in which the hidden state dimension is $2048$. We set the pruning ratio as $0.5$ for MLP and $0.5$ for Attention Layer, when pruning MLP and attention weights at the same time, we have pruning ratio $\rho$ of the whole model as $0.5$. We use $8$ sentences in the C4 training dataset [115] as calibration data, and we use $128$ sentences in the C4 validation dataset [115] to evaluate perplexity. We set $0.005$ as the learning rate $\eta$ and $0.05$ as the regularization coefficient $\lambda$.
Results. As we can see in Table 1, our Attention pruning methods always outperform Wanda Attention pruning and SparseGPT Attention pruning by a large margin when combined with Dense/Wanda/SparseGPT MLP pruning methods. These empirical results support our theoretical analysis that our pruning method can converge. Furthermore, our method is broadly applicable in the real-world case and can be combined with many other pruning methods.